OpenStudy (babyslapmafro):

Please help. University tuition and fees can be paid by using one of two plans. Early-bird: Pay total amount due 1 year in advance and get a 10% discount. On-time: Pay total amount due when classes start. The cost of tuition and fees is $10,000 per year. (a.) How much is paid in the early-bird plan? Answer: $9000 (b.) What is the equivalent amount of the savings compared to the on-time payment at the time that the on-time payment is made? Answer: ??? Please help with part (b.), thanks.

OpenStudy (anonymous):

Ok, I think it means the early-bird has to pay another 9000 by the time the on-time starts paying...

OpenStudy (anonymous):

Or maybe it just wants you to say 1000.

OpenStudy (babyslapmafro):

This is the answer: (b) Equivalent future amount = 9000(1 + 0.10) = $9900 Savings = 10,000 – 9900 = $100

OpenStudy (babyslapmafro):

I don't understand it.

OpenStudy (anonymous):

I'm so sorry I don't know either...:,(

OpenStudy (anonymous):

bump the question up.